OpenStudy (anonymous):

Using 5 as the x value, find y in the equation: -3x + 2y = 1

OpenStudy (anonymous):

Set the equation as y=...

OpenStudy (anonymous):

-3(5)+2y=1 -15+2y=1 +15 +15 2y=16 /2 /2 y=8